OpenStudy (anonymous):

Evaluate 5!. a) 60 b) 120 c) 720

OpenStudy (anonymous):

Do you know what a factorial is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It is what you are being asked to evaluate. A factorial is a product of numbers \(1\) through \(n\). \(5!=1\cdot 2\cdot 3\cdot 4\cdot 5\), for example.

OpenStudy (anonymous):

so its 120 ?

OpenStudy (anonymous):

Correct. Good job.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

You're welcome! :D